OpenStudy (anonymous):

use the formula for the area of a trapezoid. what is the area of a trapezoidal pool with a height of 15 yd and bases of 14 yd and 26 yd?

OpenStudy (anonymous):

300

OpenStudy (anonymous):

A=14+26/2 h=14+26/2*15=200